# GAS: Generative Activation-Aided Asynchronous Split Federated Learning

### v1

@@ Abstract
Split Federated Learning (SFL) splits and collaboratively trains a shared model between clients and server, where clients transmit activations and client-side models to server for updates. Recent SFL studies assume synchronous transmission of activations and client-side models from clients to server. However, due to significant variations in computational and communication capabilities among clients, activations and client-side models arrive at server asynchronously. The delay caused by asynchrony significantly degrades the performance of SFL. To address this issue, we consider an asynchronous SFL framework, where an activation buffer and a model buffer are embedded on the server to manage the asynchronously transmitted activations and client-side models, respectively. Furthermore, as asynchronous activation transmissions cause the buffer to frequently receive activations from resource-rich clients, leading to biased updates of the server-side model, we propose Generative activations-aided Asynchronous SFL (GAS). In GAS, the server maintains an activation distribution for each label based on received activations and generates activations from these distributions according to the degree of bias. These generative activations are then used to assist in updating the server-side model, ensuring more accurate updates. We derive a tighter convergence bound, and our experiments demonstrate the effectiveness of the proposed method.
@@ Introduction
Traditional SFL always assumes synchronous model exchange, where the server waits to receive all client-side models for aggregation. However, since clients have different communication and computational capabilities, client-side models are uploaded at the server asynchronously. The slow clients are referred to as stragglers, delaying the overall training process. Previous works in FL such as FedAsync (Xie, Koyejo, and Gupta 2019) and FedBuff (Nguyen et al. 2022) are proposed to tackle the stragglers issue by allowing clients to update the global model asynchronously. Additionally, CA2FL (Wang et al. 2023) addresses convergence degradation by caching and reusing previous updates for global calibration, ensuring more consistent model updates despite asynchronous conditions.
@@
  • Several useful insights are obtained via our theoretical analysis: First, the proposed GAS can mitigate the gradient dissimilarity introduced by stragglers. Second, the proposed GAS achieves a tighter convergence bound. Third, by setting a decaying learning rate, the impact of stragglers can be gradually mitigated as the training progresses.
@@
Asynchronous FL addresses the limitations of traditional synchronous FL in heterogeneous environments, where “stragglers”, or slow clients, can degrade overall training performance and efficiency (Wang et al. 2021). Early asynchronous federated learning frameworks (Xie, Koyejo, and Gupta 2019; Chen, Sun, and Jin 2019) mitigate the impact of stragglers on model performance by adaptively weighting the local updates. ASO-Fed (Chen et al. 2020) employs a dynamic learning strategy to adjust the local training step size, reducing the staleness effects caused by stragglers. FedBuff (Nguyen et al. 2022) introduces a buffering mechanism to temporarily store updates from faster clients, achieving higher concurrency and improving training efficiency. CA2FL (Wang et al. 2023) further advances this approach by caching and calibrating updates based on data properties to handle both stragglers and data heterogeneity. FedCompass (Li et al. 2023) enhances efficiency by using a computing power-aware scheduler to prioritize updates from more powerful clients, thus reducing the waiting time for stragglers. FedASMU (Liu et al. 2024) addresses the stragglers issue through dynamic model aggregation and adaptive local model adjustment methods. Moreover, some works (Lee and Lee 2021; Wang et al. 2022b; Zhu et al. 2022; Hu, Chen, and Larsson 2023) have further optimizes the performance of asynchronous FL in wireless communication environments through staleness-aware model aggregation schemes and client selection strategies.
Note that the asynchronous transmissions of activations and client-side models in SFL present unique challenges that prevent the direct application of previous asynchronous FL methods. Furthermore, the more frequent transmissions of activations exacerbate the stragglers issue. Existing SFL methods (Yan et al. 2023; Shen et al. 2023) employ adaptive model splitting to balance activation arrival times to mitigate the stragglers issue. However, these methods are constrained by the model structure. The above challenges highlight the need for further research to develop a tailored asynchronous framework for SFL. In this paper, we propose GAS to fill the gap by introducing a novel buffer mechanism and generative activations, which address the stragglers issue in SFL and achieve better model performance and training efficiency.
@@
For the baseline comparison, we include both asynchronous and synchronous FL algorithms. The baseline asynchronous FL algorithms are FedBuff (Nguyen et al. 2022) and CA2FL (Wang et al. 2023). FedBuff introduces a buffer mechanism to enable asynchronous FL, while CA2FL builds on FedBuff by incorporating cached update calibration to enhance model performance in the presence of client data heterogeneity. Additionally, we select MiniBatch-SFL (Huang, Tian, and Tang 2023) and S2FL (Yan et al. 2023) as baseline synchronous SFL algorithms. MiniBatch-SFL improves SFL performance by updating server-side model centrally, while S2FL builds on MiniBatch-SFL by introducing adaptive model splitting and activation grouping strategies to address the stragglers issue.
@@ E. More Experimental Details
We conducted our experiments using two RTX 4080 GPUs. For all datasets, we used the SGD optimizer with a learning rate of $0.01$, momentum of $0.9$, and weight decay of $0.0005$. For each training task, we ran experiments with random seeds of $2023$, $1998$, and $1125$. Full details are provided in the supplementary materials along with the code.
@@ G. Ablation Study on Splitting layer
H. Enhancing Privacy in GAS

### v2

@@ Abstract
Split Federated Learning (SFL) splits and collaboratively trains a shared model between clients and server, where clients transmit activations and client-side models to server for updates. Recent SFL studies assume synchronous transmission of activations and client-side models from clients to server. However, due to significant variations in computational and communication capabilities among clients, activations and client-side models arrive at server asynchronously. The delay caused by asynchrony significantly degrades the performance of SFL. To address this issue, we consider an asynchronous SFL framework, where an activation buffer and a model buffer are embedded on the server to manage the asynchronously transmitted activations and client-side models, respectively. Furthermore, as asynchronous activation transmissions cause the buffer to frequently receive activations from resource-rich clients, leading to biased updates of the server-side model, we propose Generative activations-aided Asynchronous SFL (GAS). In GAS, the server maintains an activation distribution for each label based on received activations and generates activations from these distributions according to the degree of bias. These generative activations are then used to assist in updating the server-side model, ensuring more accurate updates. We derive a tighter convergence bound, and our experiments demonstrate the effectiveness of the proposed method. The code is available at https://github.com/eejiarong/GAS.
@@ Introduction
Traditional SFL always assumes synchronous model exchange, where the server waits to receive all client-side models for aggregation. However, since clients have different communication and computational capabilities, client-side models are uploaded at the server asynchronously. The slow clients are referred to as stragglers, delaying the overall training process. Previous works in FL such as FedAsync (Xie, Koyejo, and Gupta 2019) and FedBuff (Nguyen et al. 2022) are proposed to tackle the stragglers issue by allowing clients to update the global model asynchronously. Additionally, CA2FL (Wang et al. 2024) addresses convergence degradation by caching and reusing previous updates for global calibration, ensuring more consistent model updates despite asynchronous conditions.
@@
  • Several useful insights are obtained via our theoretical analysis: First, GAS can mitigate the gradient dissimilarity introduced by stragglers. Second, GAS achieves a tighter convergence bound. Third, by setting a decaying learning rate, the impact of stragglers can be gradually mitigated as the training progresses.
@@
Asynchronous FL addresses the limitations of traditional synchronous FL in heterogeneous environments, where “stragglers”, or slow clients, can degrade overall training performance and efficiency (Wang et al. 2021). Early asynchronous FL frameworks (Xie, Koyejo, and Gupta 2019; Chen, Sun, and Jin 2019) mitigate the impact of stragglers by adaptively weighting the local updates. ASO-Fed (Chen et al. 2020) employs a dynamic learning strategy to adjust the local training step size, reducing the staleness effects caused by stragglers. FedBuff (Nguyen et al. 2022) introduces a buffering mechanism to temporarily store updates from faster clients, achieving higher concurrency and improving training efficiency. CA2FL (Wang et al. 2024) further advances this approach by caching and calibrating updates based on data properties to handle both stragglers and data heterogeneity. FedCompass (Li et al. 2023) enhances efficiency by using a computing power-aware scheduler to prioritize updates from more powerful clients, thus reducing the waiting time for stragglers. FedASMU (Liu et al. 2024) addresses the stragglers issue through dynamic model aggregation and adaptive local model adjustment methods. Moreover, some works (Lee and Lee 2021; Wang et al. 2022b; Zhu et al. 2022; Hu, Chen, and Larsson 2023) have further optimizes the performance of asynchronous FL in wireless communication environments through staleness-aware model aggregation and client selection schemes.
Note that the asynchronously transmitted activations are concatenated rather than aggregated to update the server-side model in SFL, introducing unique challenges that make previous asynchronous FL methods inapplicable. Furthermore, the more frequent transmissions of activations exacerbate the stragglers issue. Existing SFL methods (Yan et al. 2023; Shen et al. 2023) employ adaptive model splitting to balance activation arrival times; however, they are constrained by the model structure. The above challenges highlight the need for further research to develop a tailored asynchronous framework for SFL. In this paper, we propose GAS to fill the gap by introducing a novel buffer mechanism and generative activations, which address the stragglers issue in SFL and achieve better model performance. Additionally, while CCVR (Luo et al. 2021) and FedImpro (Tang et al. 2024) also employ activation generation to enhance model performance, they require the additional transmission of local activation distributions. GAS distinguishes itself by leveraging the inherent characteristics of the SFL framework to dynamically update activation distributions using the activations continuously uploaded by clients, without incurring extra communication overhead.
@@
For the baseline comparison, we include both asynchronous and synchronous FL algorithms. The baseline asynchronous FL algorithms are FedBuff (Nguyen et al. 2022) and CA2FL (Wang et al. 2024). FedBuff introduces a buffer mechanism to enable asynchronous FL, while CA2FL builds on FedBuff by incorporating cached update calibration to enhance model performance in the presence of client data heterogeneity. Additionally, we select MiniBatch-SFL (Huang, Tian, and Tang 2023) and S2FL (Yan et al. 2023) as baseline synchronous SFL algorithms. MiniBatch-SFL improves SFL performance by updating server-side model centrally, while S2FL builds on MiniBatch-SFL by introducing adaptive model splitting and activation grouping strategies to address the stragglers issue.
@@ E. More Experimental Details
We conducted our experiments using two RTX 4080 GPUs. For all datasets, we used the SGD optimizer with a learning rate of $0.01$, momentum of $0.9$, and weight decay of $0.0005$. For each training task, we ran experiments with random seeds of $2023$, $1998$, and $1125$. Full details are provided in the code.
@@ G. Ablation Study on Splitting layer
Method CIFAR-10
$\text{s}=2$ $\alpha=0.1$
ViT-small FedAvg $70.85_{\pm 0.68}$ $68.95_{\pm 1.36}$
FedBuff $58.46_{\pm 2.18}$ $59.84_{\pm 3.64}$
CA2FL $68.27_{\pm 0.11}$ $70.58_{\pm 1.20}$
Ours $\mathbf{79.81}_{\pm 1.43}$ $\mathbf{72.89}_{\pm 4.75}$
ResNet18 FedAvg $34.47_{\pm 2.51}$ $75.39_{\pm 7.04}$
FedBuff $31.83_{\pm 3.45}$ $72.73_{\pm 6.71}$
CA2FL $37.81_{\pm 8.77}$ $76.28_{\pm 7.16}$
Ours $\mathbf{83.58}_{\pm 5.62}$ $\mathbf{80.73}_{\pm 1.21}$
Table 5: Test accuracy (%) under different model structures.
H. Ablation Study on Model Structures
We conduct supplementary experiments using ViT-small (Han et al. 2022) and ResNet18 (He et al. 2016). In the ViT-small experiments, we employ a compact Vision Transformer architecture consisting of $6$ Transformer layers, each with $8$ attention heads and a hidden MLP dimension of $512$. We set the dropout rate to $0.1$ for both the Transformer layers and the embedding layer. The split point of the model is set after the first Transformer layer, with a learning rate of $0.0001$, optimized using the Adam optimizer. In the ResNet18 supplementary experiments, the model is split after the first residual layer, and we use a learning rate of $0.005$ with the SGD optimizer. The experimental results, as shown in the Table 5, demonstrate that the proposed GAS exhibits good scalability and outperforms benchmarks across various levels of data heterogeneity.
I. Enhancing Privacy in GAS